OpenStudy (anonymous):

-4y=-64 - 6x 4x - 4y=-36

OpenStudy (anonymous):

ok so this one should be the same thing we've been doing. add 6x to the other side in the first equation to get 6x-4y=-64. multiply this equation by -1 to get -6x+4y=64. add the two equations up to get -2x=28 so x=-14. substitute this into the first or second equation. I'll do it to the second equation this time. 4(-14)-4y=-36. -56-4y=-36. -4y=20. so y=5.

OpenStudy (anonymous):

whew the questions keep getting longer and longer. and i'm getting reaaaalllly tired. so hard to do it online lol..

OpenStudy (anonymous):

x=-14

OpenStudy (anonymous):

i said that.

OpenStudy (anonymous):

kk